OpenStudy (anonymous):

Using set builder notation, what is the domain for the function f(x) = 1/x+1 ?

OpenStudy (anonymous):

\[f(x) = \frac{ 1 }{ x+1 }\]

jimthompson5910 (jim_thompson5910):

what value of x makes the denominator equal to zero?

OpenStudy (anonymous):

-1 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you get that by setting the denominator x+1 equal to zero and solving for x good

jimthompson5910 (jim_thompson5910):

this x value makes the denominator zero, so we have to exclude it out of the domain every other number is allowed because any other number doesn't make the denominator zero

jimthompson5910 (jim_thompson5910):

using set builder notation, the domain is... \[\Large D = \left\{ x|x\in\mathbb{R}, x\neq -1 \right\}\]

jimthompson5910 (jim_thompson5910):

this basically says: the domain is the set x such that x is any real number but x can't be -1

OpenStudy (anonymous):

ohhhh okay thanks

jimthompson5910 (jim_thompson5910):

you're welcome